OpenStudy (rane):

james was driving his car on a level road, in a 60km/h zone at a constant speed, when he accidentally hit a low rock wall. Fortunately, James was wearing his seat belt and was unhurt in the accident. Unfortunately, Taymen was car-suffering at the time and, on impact, came off the roof, striking the ground 12m in front f the point where he was standing on the car's roof.

OpenStudy (rane):

Tianyan, the investigation police officer, was called to the scene and, after taking some measurements, and making certain assumptions, came up with an estimate of the speed of James car when it collided with the wall. List her assumptions and determine the estimated speed. the height of the car = 1.5m the distance between the Taymen came off the roof to the point where he/she landed = 12m

OpenStudy (rane):

these are the some measurements that pounce officer took

OpenStudy (anonymous):

do u have any idea urself on how to solve it??????

OpenStudy (rane):

nop… i tried but i got it wrong so i don't know

OpenStudy (anonymous):

ok.now u see that the car was having velocity only in the horizontal direction.when the car stopped,the man had velocity only in the horizontal direction.so we shud first solve for vertical direction because the man's initial velocity in vertical direction is zero,u have the vertical distance travelled and the acceleration is 9.8 due to gravity.now with this data try to find the time taken by man to fall on the ground.

OpenStudy (anonymous):

did u get it???????

OpenStudy (rane):

\[t = \sqrt{\frac{ 2*1.5 }{ 9.8}} = 0\]

OpenStudy (rane):

0.55s

OpenStudy (anonymous):

yes.so now u have the total time.now solving for the horizontal direction, u have time and the distance and the velocity was constant in the horizontal direction.so u can find the horizontal velocity of the man which is the same as of the car.

OpenStudy (rane):

\[v = \frac{ 12 }{ 0.55 } = 21.8 m/s \]

OpenStudy (anonymous):

its correct

OpenStudy (rane):

can u help me with one more ?

OpenStudy (anonymous):

ya.sure

OpenStudy (rane):

using your result from above, and assuming that James & Taymen have similar masses, estimate the impulse for each oh them arising frm the accident and compare the two results

OpenStudy (anonymous):

impulse=change in momentum as u can see taymen's velocity didnt change,his motion continued after accident with same velocity,so for him no change in momentum. but james suddenly stopped, so there was a change in his momentum. so for james, impulse=final momentum-initial momentum

OpenStudy (anonymous):

the negative sign in this case will show the direction of impulse

OpenStudy (rane):

sp i don't have to find the momentum just the explanation ?

OpenStudy (rane):

so*

OpenStudy (anonymous):

for taymen it is zero as his velocity is same before and after the accident for james u need to find it

OpenStudy (rane):

wht abt the masses?

OpenStudy (anonymous):

didnt thought about that.so i guess u shud just go wid the explanation

OpenStudy (rane):

yh thats wht i thought as well